OpenStudy (anonymous):

Use implicit differentiation to find dy/dx 2xy - y^2 = 1 2xy - 2y * y' = 1 ? How do I do this problem?

terenzreignz (terenzreignz):

The gist of it is... you differentiate x like normal and you differentiate y as you would for x, except after you're done differentiating the y-part, you multiply it to dy/dx or y' For instance, the derivative of x^3 is still 3x^2 but the derivative of y^3 is (3y^2)(dy/dx) It's like the chain rule, see..

OpenStudy (anonymous):

okay well so far I have 2y + 2xy' - 2y * y' = 1 Not sure if that's right.

terenzreignz (terenzreignz):

You forgot to differentiate the right-side of the equation. Otherwise, you've done well :P

OpenStudy (anonymous):

What would that be? I was going to leave the one there and move my x's and y's to the right.

terenzreignz (terenzreignz):

If it were 2xy - y^2 = x, then certainly, differentiating gives 2y + 2xy' - (2y)(y') = 1 But it's 1 on the other side, and you differentiated the left...

terenzreignz (terenzreignz):

In general, when you have two functions... such that f(x) = g(x) You can take the derivative of one side, but are you SURE it's still equal to the other function? Most of the time it isn't, BUT it's true that f'(x) = g'(x)

OpenStudy (raden):

derivative of 1 is 0